OpenStudy (anonymous):

Solve x2 + 10x + 7 = 0 by completing the square. Which equation is used in the process?

OpenStudy (anonymous):

(x + 10)^2 = 93 (x + 5)^2 = 32 (x + 5)^2 = 18

OpenStudy (anonymous):

(x+5)^2=32

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[x ^{2}+10x+(10/2)^{2}=-7+(10/2)^{2}\] and then you have \[(x+5)^{2}= 18\] Made a mistake previously.